# RsGCN: Rescaling Enhances Generalization of GCNs for Solving Scalable Traveling Salesman Problems

### v3

@@ Appendix I Limitations and Broader Impacts
Our current methodology focuses on solving Symmetric TSP. Future work requires further extensions to broader routing problems, including the Asymmetric TSP (ATSP) and Capacitated Vehicle Routing Problem (CVRP). Our proposed solver still trails behind Concorde since Concorde is an exact solver that can obtain the optimal tour of almost every instance. However, Concorde highly relies on expert-crafted heuristics specifically designed for TSPs. LKH is another solver that also highly relies on expert-crafted heuristics, while the proposed solver can achieve comparable performance to LKH on large-scale TSPs with more than 1K nodes. In addition, the proposed solver significantly outperforms neural baselines.
@@
Our work highlights the importance of rescaling scale-dependent features and provides insightful implications for future research to improve the generalization of neural combinatorial optimization (NCO) solvers. Furthermore, our solver’s small parameter scale and low training cost will prompt the community to rethink the practical utility and cost-effectiveness of NCO solvers.

### v4

@@ Appendix H Limitations and Broader Impacts
Our current methodology focuses on solving Symmetric TSP. Future work requires further extensions to broader routing problems, including the Asymmetric TSP (ATSP) and Capacitated Vehicle Routing Problem (CVRP). Our proposed methods still trails behind Concorde and LKH-3. However, Concorde and LKH-3 highly rely on expert-crafted heuristics specifically designed for TSPs.
@@
Our work highlights the importance of subgraph-based rescaling and provides insightful implications for future research to improve the generalization of neural combinatorial optimization (NCO) solvers. Furthermore, our solver’s small parameter scale and low training cost will prompt the community to rethink the practical utility and cost-effectiveness of NCO solvers.
Figure 11: Visualization of the subgraph-based rescaling on TSP-20/200/2K.